ankit is now 9 yrs older thn bobby.in 10 yrs ankit will be twice as old as bobby was 10 yrs ago.find their age i tried this sum plenty of times but i didnt get it right please help me and thanks


Answer by ankor@dixie-net.com(22740)

Ankit is now 9 yrs older than bobby; in 10 yrs ankit will be twice as old as bobby was 10 yrs ago. find their ages
:
Write an expression for each statement
"Ankit is now 9 yrs older than bobby"
B = Bobby's age now
Then
A = (B + 9) (Ankit's age)
:
"in 10 yrs ankit will be twice as old as bobby was 10 yrs ago."
(A + 10) = 2(B - 10)
A + 10 = 2B - 20
A = 2B - 20 - 10
A = 2B - 30
:
Substitute (B+9) for A in the above equation; solve for B:
B + 9 = 2B - 30
B = 2B - 30 - 9
B - 2B = -39
-B = -39
or
B = +39 is Bobby's age now
Then
39 + 9 = 48 = Ank's now
:
:
Check our solution in the statememt:
"in 10 yrs ankit will be twice as old as bobby was 10 yrs ago"
(48 + 10) = 2(39 - 10)
58 = 2(29) confirms or solutions
:
Did this method help with this stuff?